OpenStudy (anonymous):

I need help figuring out how to get a certain number of packets that equal a specific price please help! Vanilla wafers are $4 and Chocolate is $1. If I spent $45 on 27 packets of wafers how many of both would I have of each?

OpenStudy (texaschic101):

if you mean vanilla wafers and chocolate wafers.... v + c = 27 4v + 1c = 45 now you can solve by substitution

OpenStudy (anonymous):

I have tried that and I keep messing up and I just cant figure it out.

TheSmartOne (thesmartone):

How have you tried that?

OpenStudy (texaschic101):

v + c = 27 v = 27 - c 4v + 1c = 45 4(27 - c) + 1c = 45 108 - 4c + 1c = 45 -4c + 1c = 45 - 108 - 3c = - 63 c = 21 v + c = 27 v + 21 = 27 v = 27 - 21 v = 6

OpenStudy (anonymous):

because I have used that already

OpenStudy (anonymous):

@texaschic101 thanks so much! :)

OpenStudy (texaschic101):

did you find where you messed up ?